Please help me solve these two equations:
1. y = 2 - 6x
1/2y - x = 1
2. 1/x + 1/y = 1
3/x - 2/y = 8

Answer by stanbon(75887)

y = 2 - 6x
(1/2)y - x = 1
Sustitute for "y" and solve for "x":
(1/2)(2-6x) -x = 1
1-3x -x = 1
4x = 0
x = 0
----
Solve for "y":
y = 2 - 6x
y = 2 - 6*0
y = 2
=================

1/x + 1/y = 1
3/x - 2/y = 8
------
Multiply thru the 1st equation by 2:
2/x + 2/y = 2
3/x - 2/y = 8
-----
Add to get:
5/x = 10
x = 1/2
-----
Solve for "y":
1/x + 1/y = 1
2 + 1/y = 1
1/y = -1
y = -1
======================
Cheers,
Stan H.

-----------------------